OpenStudy (anonymous):

The strength of an electric field at 5.0 cm from a point charge is 100.0 N/C. What is the magnitude of the source charge? Show your work. Could anyone help me understand this? Physics is not my strongest subject.

OpenStudy (anonymous):

let the source charge be q and its electric field be E which is given as 100 N/C Consider a point P from the charge at a distance r which is 5 cm. r = 5/100 = 0.05 m (because all the units should be according to the si units) the electric field due to a point charge is given by \[E = \frac{ q }{ 4\pi \epsilon * r ^{2} } \] (derived from coulomb's law) I have taken for granted that the system is placed in air so the value of the constant is 9 * 10^9 \[100 = \frac{ 9 * 10^{9} * q }{ 0.05 }\] \[q = 0.55 * 10^{-9} C\]